Sammy takes 5 hours longer to complet 300 lines of code than Martin. Together Sammy and Martin can complete the code in 6 hours.
How long it takes each programmer to complete the code if they program on their own. (hint: Quadratic equaiton )and 2) at what speed they are programming lines per hour.

Answer by checkley77(12844)

1/X+1/(X+5)=1/6
(X+5+X)/X(X+5)=1/6
(2X+5)/(X^2+5X)=1/6 CROSS MULTIPLY
X^2+5X=6(2X+5)
X^2+5X=12X+30
X^2+5X-12X-30=0
X^2-7X-30=0
(X-10)(X+3)=0
X-10=0
X=10 HOURS FOR MARTIN CODING ALONE.
10+5=15 HOURS FOR SAMMY CODING ALONE.
Proof:
1/10+1/15=1/6
(15+10)/150=1/6
25/150=1/6
1/6=1/6
300/10=30 LINES OF CODE PER HOUR FOR MARTIN.
300/15=20 LINES OF CODE PER HOUR FOR SAMMY.